# Thresholded Lasso for high dimensional variable selection

### v2

@@ Appendix A Proofs
It holds by definition of $S_{{\mathcal{D}}}$ that $I\cap S_{{\mathcal{D}}}=\emptyset$. One can check via the proof of Theorem 2.4 that (19) holds for $D_{0}^{\prime},D_{1}$ as defined in (2.1) and (26) respectively. Hence by Lemma 2.3, we have on event ${\mathcal{T}}_{\alpha}$ for $C_{4}\geq D_{1}$, $|I|\leq 2s_{0}$ and $|I\cup S_{{\mathcal{D}}}|\leq|I\cup S|\leq s+s_{0}\leq 2s$, given that $|{S_{{\mathcal{D}}}}|<s$ and moreover
$\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}$$\leq$ $\sqrt{(D^{\prime}_{0}+C_{4})^{2}+1}\lambda\sigma\sqrt{s_{0}}.$
We have by Lemma 2.7, on event ${\mathcal{T}}_{\alpha}$, for $\lambda=\sqrt{2\log p/n}$, $\left\lvert I\right\rvert<2s_{0}$, for $\hat{\beta}=\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$
$\left\lVert\hat{\beta}-\beta\right\rVert_{2}^{2}:=\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}$
@@
$\leq$ $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\left(1+\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda_{\min}^{2}(|I|)}\right)+\frac{2|I|(1+a)\sigma^{2}\lambda^{2}}{\Lambda^{2}_{\min}(|I|)}$
@@
$\leq$ $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\left(1+\frac{2\theta^{2}_{\left\lvert S_{{\mathcal{D}}}\right\rvert,|I|}}{\Lambda_{\min}^{2}(|I|)}\right)+\frac{4(1+a)}{\Lambda_{\min}^{2}(2s_{0})}\sigma^{2}\lambda^{2}s_{0}$
@@
$\leq$ $\lambda^{2}\sigma^{2}s_{0}((D^{\prime}_{0}+C_{4})^{2}+1)\left(1+\frac{2\theta^{2}_{\left\lvert S_{{\mathcal{D}}}\right\rvert,|I|}}{\Lambda_{\min}^{2}(|I|)}+\frac{4}{9}\right),$
@@
where we use the fact that ${4(1+a)}/{\Lambda_{\min}^{2}(2s_{0})}\leq\frac{4}{9}(D^{\prime}_{0})^{2}$ since
@@
$K^{4}(s_{0},4)$$\geq$ $K^{4}(s_{0},1)\geq 1/{(4\Lambda^{2}_{\min}(2s_{0}))}$
in view of (35). Now (17) clearly holds with
$D_{4}^{2}$ $=$ $((D^{\prime}_{0}+C_{4})^{2}+1)\left(1+\frac{2\theta_{|I|,|{S_{{\mathcal{D}}}}|}^{2}}{\Lambda_{\min}^{2}(2s_{0})}+\frac{4}{9}\right).$
Now (18) holds by Lemma 2.8. ∎
A.2 Proof of Proposition 4.1
Now given that $\beta_{i}\geq\beta_{j}$ for all $i\in T_{0},j\in T_{0}^{c}$, the proposition holds. ∎
A.3 Proof of Lemma 2.3
@@
Let $\beta_{{\mathcal{D}}}=\beta_{{\mathcal{D}}}^{(1)}+\beta_{{\mathcal{D}}}^{(2)}$, where $\beta_{{\mathcal{D}}}^{(1)}:=(\beta_{j})_{j\in T_{0}\cap{\mathcal{D}}}$ consists of coefficients that are significant relative to $\lambda\sigma$, but are dropped as $\beta_{j,\text{\rm init}}<t_{0}$, and $\beta_{{\mathcal{D}}}^{(2)}$ consists of those below $\lambda\sigma$ in magnitude that are dropped. Hence
@@
where the second equality is by (14) and the last inequality is by definition of $s_{0}$ as in (13). Now for $\beta_{{\mathcal{D}}}^{(1)}$, where $|{\mathcal{D}}_{1}|<s_{0}$, we have by the triangle inequality,
$\left\lVert\beta_{{\mathcal{D}}}^{(1)}\right\rVert_{2}$$\leq$ $\left\lVert\beta_{{\mathcal{D}}_{1},\text{\rm init}}\right\rVert_{2}+\left\lVert\beta_{{\mathcal{D}}_{1},\text{\rm init}}-\beta_{{\mathcal{D}}}^{(1)}\right\rVert_{2}$ (92)
@@
where we used the fact that $\left\lVert\beta_{{\mathcal{D}}_{1},\text{\rm init}}-\beta_{{\mathcal{D}}}^{(1)}\right\rVert_{2}:=\left\lVert h_{{\mathcal{D}}_{1}}\right\rVert_{2}\leq\left\lVert h_{T_{0}}\right\rVert_{2}\leq D^{\prime}_{0}\lambda\sigma\sqrt{s_{0}}$ by (19). Hence (21) holds given (90), (92), and (91). ∎
Appendix B Proof of Lemmas 2.5 and 2.7
Proof of Lemma 2.5. Recall that the random variable $\left\lVert Q\right\rVert_{2}^{2}\sim\chi_{m}^{2}$ is distributed according to the chi-square distribution where $\left\lVert Q\right\rVert_{2}^{2}=\sum_{i=1}^{m}Q_{i}^{2}$ with $Q_{i}\sim N(0,1)$ that are independent and normally distributed. By [20],
@@
$=$ ${\mathbb{P}}\left(\frac{\chi^{2}_{m}}{m}-1\geq\delta\right)+{\mathbb{P}}\left(\frac{\chi^{2}_{m}}{m}-1\leq-\delta\right)$
$\leq$ $\exp(-3m\delta^{2}/(16))+\exp(-m\delta^{2}/4).$
@@
where we used an upper bound on $\left\lVert Q\right\rVert_{2}^{2}\leq 3m/2$, $\left\lVert Q\right\rVert_{2}^{2}\sim\chi^{2}_{m}$, and the fact that
@@
We now focus on bounding the first term in (95). Let $P_{I}$ denote the orthogonal projection onto $I$. Clearly, $I\cap S_{{\mathcal{D}}}=\emptyset$. Let
@@
[…]
@@
[…]
@@
Now the lemma holds given that $\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}=\hat{\beta}_{I}$, $\hat{\beta}_{I^{c}}^{\mathop{\text{\rm ols}\kern 0.60275pt}}=0$, $\beta_{I}-\beta=-\beta_{I^{c}}$, and hence
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}-\beta\right\rVert_{2}^{2}$ $=$ $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I^{c}}-\beta_{I^{c}}\right\rVert_{2}^{2}$
$=$ $\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta_{I}-\beta\right\rVert_{2}^{2}$
$\leq$ $\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda^{2}_{\min}(|I|)}\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}+\frac{3|I|\sigma^{2}}{n\Lambda_{\min}(|I|)}+\left\lVert\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}^{2},$
where $\beta_{I}$ is the the restriction of $\beta$ to the set $I$. ∎
Remark B.1.
Notice that we can also derive the lower bound on event $\mathcal{Q}^{c}$ for $\left|I\right|=m$
@@
B.1 Proof of Lemma 2.7
The only part we make modification is the following: On event ${\mathcal{T}}_{\alpha}$, we have by (2.6), for an arbitrary set $I$ of size $\left|I\right|\leq 2s$,
@@
and thus (32) holds since $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}=\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta_{I}-\beta\right\rVert_{2}^{2}$ by (B), where $\beta_{I}$ is the the restriction of $\beta$ to the set $I$. ∎
B.2 Proof of Lemma 4.2
@@
By definition of $s_{0}$ as in (13), we have $\sum_{i=1}^{p}\min(\beta_{i}^{2},\lambda^{2}\sigma^{2})\leq s_{0}\lambda^{2}\sigma^{2}$. Now it is clear that
@@
It is clear for ${\mathcal{D}}_{11}={\mathcal{D}}\cap A_{0}$, we have ${\mathcal{D}}_{11}\subset A_{0}\subset T_{0}\subset S$. Let $\beta_{{\mathcal{D}}}^{(11)}:=(\beta_{j})_{j\in A_{0}\cap{\mathcal{D}}}$. Now by (103), we have
$\;\;\;\;\;\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\leq\left\lVert\beta_{{\mathcal{D}}}^{(11)}\right\rVert_{2}^{2}+\left\lVert\beta^{(12)}+\beta^{(2)}\right\rVert_{2}^{2}\leq\left\lVert\beta_{{\mathcal{D}}}^{(11)}\right\rVert_{2}^{2}+(s_{0}-a_{0})\lambda^{2}\sigma^{2},$
where $|{\mathcal{D}}_{11}|\leq a_{0}$, $\left\lVert\beta_{{\mathcal{D}}}^{(11)}\right\rVert_{\infty}<t_{0}$ and we have by the triangle inequality,
$\left\lVert\beta_{{\mathcal{D}}}^{(11)}\right\rVert_{2}$$\leq$ $\left\lVert\beta_{{\mathcal{D}}_{11},\text{\rm init}}\right\rVert_{2}+\left\lVert\beta_{{\mathcal{D}}_{11},\text{\rm init}}-\beta_{{\mathcal{D}}}^{(11)}\right\rVert_{2}\leq t_{0}\sqrt{\left|{\mathcal{D}}_{11}\right|}+\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}$ (104)
$\leq$ $t_{0}\sqrt{a_{0}}+\left\lVert h_{{\mathcal{D}}{11}}\right\rVert_{2}.$
Thus (59) holds. Now we replace the crude bound of $|{\mathcal{D}}_{11}|\leq a_{0}$ with
$|{\mathcal{D}}_{11}|\leq\frac{\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}^{2}}{|\beta_{\min,A_{0}}-t_{0}|^{2}}$
in (104) to obtain
$\left\lVert\beta_{{\mathcal{D}}}^{(11)}\right\rVert_{2}\leq t_{0}\frac{\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}}{\beta_{\min,A_{0}}-t_{0}}+\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}=\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}\frac{\beta_{\min,A_{0}}}{\beta_{\min,A_{0}}-t_{0}},$
which proves (60). ∎
B.3 Proof of Lemma 4.3
Suppose ${\mathcal{T}}_{a}\cap Q_{c}$ holds. It is clear by the choice of $t_{0}$ in (4.3) and by (61) that
@@
Thus by (4.3), we can bound $|I\cap T_{0}^{c}|$, depending on which one is applicable, by either
$|I\cap T_{0}^{c}|\leq{\left\lVert\beta_{T_{0}^{c},\text{\rm init}}\right\rVert_{1}}/{t_{0}}\leq\breve{s}_{0}\quad\text{ or }\quad|I\cap T_{0}^{c}|\leq{\left\lVert\beta_{T_{0}^{c},\text{\rm init}}\right\rVert_{2}^{2}}/{t_{0}^{2}}\leq\breve{s}_{0}.$
Moreover, we have by Lemma 2.7, for $s_{0}\leq\breve{s}_{0}\leq s$, on event ${\mathcal{T}}_{a}$,
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}$$\leq$ $\left(\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda_{\min}^{2}(|I|)}+1\right)\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}+\frac{2(1+a)|I|\lambda^{2}\sigma^{2}}{\Lambda^{2}_{\min}(|I|)}$
$\leq$ $\left(2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}+{\Lambda_{\min}^{2}(|I|)}+4(1+a)\right)\breve{s}_{0}\lambda^{2}\sigma^{2}/\Lambda^{2}_{\min}(|I|),$
where $\left|I\right|\leq s_{0}+\breve{s}_{0}$, and $\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}$ is bounded in Lemma 2.8 given $|I|+|{S_{{\mathcal{D}}}}|\leq s+|I\cap T_{0}^{c}|\leq s+\breve{s}_{0}\leq 2s$. ∎
Appendix C Proof sketch of Theorem 3 in [43]

### v3

@@ 6 Proofs
It holds by definition of $S_{{\mathcal{D}}}$ that $I\cap S_{{\mathcal{D}}}=\emptyset$. One can check via the proof of Theorem 2.4 that (19) holds for $D_{0}^{\prime},D_{1}$ as defined in (26) and (27) respectively. Hence by Lemma 2.3, we have on event ${\mathcal{T}}_{\alpha}$ $|I|\leq 2s_{0}$ and $|I\cup S_{{\mathcal{D}}}|\leq s+s_{0}\leq 2s$ for $C_{4}\geq D_{1}$; See also (90). Moreover, $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}\leq\sqrt{(D^{\prime}_{0}+C_{4})^{2}+1}\lambda\sigma\sqrt{s_{0}}.$ We have by Lemma 2.7, on event ${\mathcal{T}}_{\alpha}$, for $\lambda=\sqrt{2\log p/n}$, $\left\lvert I\right\rvert<2s_{0}$, for $\hat{\beta}=\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$,
@@
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}$ $\leq$ $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\left(1+\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda_{\min}^{2}(|I|)}\right)+\frac{2|I|(1+a)\sigma^{2}\lambda^{2}}{\Lambda^{2}_{\min}(|I|)}$
@@
$\leq$ $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}\big(1+\frac{2\theta^{2}_{\left\lvert S_{{\mathcal{D}}}\right\rvert,|I|}}{\Lambda_{\min}^{2}(2s_{0})}\big)+\frac{4(1+a)}{\Lambda_{\min}^{2}(2s_{0})}\sigma^{2}\lambda^{2}s_{0}$
@@
$\leq$ $\lambda^{2}\sigma^{2}s_{0}((D^{\prime}_{0}+C_{4})^{2}+1)\big(1+\frac{2\theta^{2}_{\left\lvert S_{{\mathcal{D}}}\right\rvert,|I|}}{\Lambda_{\min}^{2}(2s_{0})}+\frac{4}{9}\big),$
@@
where $\Lambda_{\min}(|I|)\geq\Lambda_{\min}(2s_{0})$ by (31) since $\left\lvert I\right\rvert<2s_{0}$, and the fact that ${4(1+a)}/{\Lambda_{\min}^{2}(2s_{0})}\leq\frac{4}{9}(D^{\prime}_{0})^{2}$ since
@@
[…]
Now (17) clearly holds with $D_{4}^{2}=((D^{\prime}_{0}+C_{4})^{2}+1)\big(1+\frac{2\theta_{|I|,|{S_{{\mathcal{D}}}}|}^{2}}{\Lambda_{\min}^{2}(2s_{0})}+\frac{4}{9}\big)$, and (18) holds by Lemma 2.8. ∎
6.2 Proof of Lemma 2.3
@@
Let $\beta_{{\mathcal{D}}}=\beta_{{\mathcal{D}}}^{(1)}+\beta_{{\mathcal{D}}}^{(2)}$, where $\beta_{{\mathcal{D}}}^{(1)}:=(\beta_{j})_{j\in T_{0}\cap{\mathcal{D}}}$ consists of coefficients that are significant relative to $\lambda\sigma$, but are dropped as $\left\lvert\beta_{j,\text{\rm init}}\right\rvert<t_{0}$, and $\beta_{{\mathcal{D}}}^{(2)}$ consists of those strictly below $\lambda\sigma$ in magnitude that are dropped; cf. (14). Hence
@@
where the second equality is by (14) and the last inequality is by definition of $s_{0}$ as in (13). Let ${\mathcal{D}}_{1}={\mathcal{D}}\cap T_{0}$, where $|{\mathcal{D}}_{1}|<s_{0}$. Now by the triangle inequality,
$\left\lVert\beta_{{\mathcal{D}}}^{(1)}\right\rVert_{2}=\left\lVert\beta_{{\mathcal{D}}_{1}}\right\rVert_{2}$ $\leq$ $\left\lVert\beta_{{\mathcal{D}}_{1},\text{\rm init}}\right\rVert_{2}+\left\lVert\beta_{{\mathcal{D}}_{1},\text{\rm init}}-\beta_{{\mathcal{D}}_{1}}\right\rVert_{2}$ (93)
@@
where we used the fact that $\left\lVert\beta_{{\mathcal{D}}_{1},\text{\rm init}}-\beta_{{\mathcal{D}}_{1}}\right\rVert_{2}=\left\lVert h_{{\mathcal{D}}_{1}}\right\rVert_{2}\leq\left\lVert h_{T_{0}}\right\rVert_{2}\leq D^{\prime}_{0}\lambda\sigma\sqrt{s_{0}}$ by (19). Hence (21) holds given (91), (92), and (93). ∎
6.3 Proof of Proposition 4.1
Now given that $\left\lvert\beta_{i}\right\rvert\geq\left\lvert\beta_{j}\right\rvert$ for all $i\in T_{0},j\in T_{0}^{c}$, the proposition holds. ∎
6.4 Proof of Lemma 2.5
Recall that the random variable $\left\lVert Q\right\rVert_{2}^{2}\sim\chi_{m}^{2}$ is distributed according to the chi-square distribution where $\left\lVert Q\right\rVert_{2}^{2}=\sum_{i=1}^{m}Q_{i}^{2}$ with $Q_{i}\sim N(0,1)$ that are independent and normally distributed. By [20],
@@
$=$ ${\mathbb{P}}\left(\frac{\chi^{2}_{m}}{m}-1\geq\delta\right)+{\mathbb{P}}\left(\frac{\chi^{2}_{m}}{m}-1\leq-\delta\right)\leq\exp(-3m\delta^{2}/(16))+\exp(-m\delta^{2}/4).$
@@
where we used an upper bound on $\left\lVert Q\right\rVert_{2}^{2}\leq 3m/2$, and the fact that
@@
Here $\left\lVert Q\right\rVert_{2}^{2}\sim\chi^{2}_{m}$. We now focus on bounding the first term in (96). Let $P_{I}$ denote the orthogonal projection onto $I$. Clearly, $I\cap S_{{\mathcal{D}}}=\emptyset$. Let
@@
 $\;\text{ where }\;\;\left\lVert c\right\rVert_{2}$ $\leq$ $\frac{\left\lVert X_{I}c\right\rVert_{2}}{\sqrt{n\Lambda_{\min}(|I|)}}\leq\frac{\left\lVert P_{I}X_{{S_{{\mathcal{D}}}}}\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}}{\sqrt{n\Lambda_{\min}(|I|)}}.$ (99)
@@
$\text{ and }\quad\left\lVert c\right\rVert_{2}$ $\leq$ ${\theta_{|I|,|{S_{{\mathcal{D}}}}|}\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}}/{\Lambda_{\min}(|I|)}.$ (100)
@@
Now the lemma holds given that $\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}=\hat{\beta}_{I}$, $\hat{\beta}_{I^{c}}^{\mathop{\text{\rm ols}\kern 0.60275pt}}=0$, and hence
$\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}-\beta\right\rVert_{2}^{2}$ $=$ $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta_{I^{c}}\right\rVert_{2}^{2}=\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta_{I^{c}}\right\rVert_{2}^{2}$
$\leq$ $\frac{2\theta^{2}_{|I|,|{S_{{\mathcal{D}}}}|}}{\Lambda^{2}_{\min}(|I|)}\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}+\frac{3|I|\sigma^{2}}{n\Lambda_{\min}(|I|)}+\left\lVert\beta_{{S_{{\mathcal{D}}}}}\right\rVert_{2}^{2}.$
Remark 6.1.
We can also derive the lower bound on event $\mathcal{Q}^{c}$ for $\left|I\right|=m$
@@
6.5 Proof of Lemma 2.7
The only part we make modification is the following: On event ${\mathcal{T}}_{\alpha}$, we have by (31), for an arbitrary set $I$ of size $\left|I\right|\leq 2s$,
@@
and thus (33) holds since $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}^{2}=\left\lVert\hat{\beta}_{I}-\beta_{I}\right\rVert_{2}^{2}+\left\lVert\beta_{I^{c}}\right\rVert_{2}^{2}$ by (6.4). ∎
6.6 Proof of Lemma 4.2
@@
By definition of $s_{0}$ as in (13), we have $\sum_{i=1}^{p}\min(\beta_{i}^{2},\lambda^{2}\sigma^{2})\leq s_{0}\lambda^{2}\sigma^{2}$. Now
@@
For ${\mathcal{D}}_{11}={\mathcal{D}}\cap A_{0}$, we have ${\mathcal{D}}_{11}\subset A_{0}\subset T_{0}\subset S$, and $|{\mathcal{D}}_{11}|\leq a_{0}$. Then $\left\lVert\beta_{{\mathcal{D}}_{11},\text{\rm init}}\right\rVert_{\infty}<t_{0}$. Let $\beta_{{\mathcal{D}}}^{(11)}:=(\beta_{j})_{j\in A_{0}\cap{\mathcal{D}}}$. Now by the triangle inequality,
$\left\lVert\beta_{{\mathcal{D}}}^{(11)}\right\rVert_{2}$ $=$ $\left\lVert\beta_{{\mathcal{D}}_{11}}\right\rVert_{2}\leq\left\lVert\beta_{{\mathcal{D}}_{11},\text{\rm init}}\right\rVert_{2}+\left\lVert\beta_{{\mathcal{D}}_{11},\text{\rm init}}-\beta_{{\mathcal{D}}_{11}}\right\rVert_{2}$ (105)
$\leq$ $t_{0}\sqrt{\left|{\mathcal{D}}_{11}\right|}+\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}\leq t_{0}\sqrt{a_{0}}+\left\lVert h_{{\mathcal{D}}{11}}\right\rVert_{2};$
[…]
Now (57) holds when we replace the crude bound of $|{\mathcal{D}}_{11}|\leq a_{0}$ with
[…]
$\left\lVert\beta_{{\mathcal{D}}}^{(11)}\right\rVert_{2}$ $\leq$ $t_{0}\frac{\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}}{\beta_{\min,A_{0}}-t_{0}}+\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}=\left\lVert h_{{\mathcal{D}}_{11}}\right\rVert_{2}\frac{\beta_{\min,A_{0}}}{\beta_{\min,A_{0}}-t_{0}}.$
6.7 Proof of Lemma 4.3
By the choice of $t_{0}$ in (4.3) and by (58),
@@
Thus by (4.3), we can bound $|I\cap T_{0}^{c}|$, by either
$|I\cap T_{0}^{c}|\leq{\left\lVert\beta_{T_{0}^{c},\text{\rm init}}\right\rVert_{1}}/{t_{0}}\leq\breve{s}_{0}\quad\text{ or }\quad|I\cap T_{0}^{c}|\leq{\left\lVert\beta_{T_{0}^{c},\text{\rm init}}\right\rVert_{2}^{2}}/{t_{0}^{2}}\leq\breve{s}_{0},$
depending on which one is applicable. The rest follows from the proof of Lemma 4.2. ∎
7 Proof sketch of Theorem 3 in [43]